### v1

# Technical Note: Vendor-Specific Approach for Standardized Uptake Value Calculation

Abstract
Background: The Standardized Uptake Value (SUV) is a critical metric in positron emission tomography (PET) imaging, used to assess metabolic activity. However, calculating SUV from DICOM files presents challenges due to vendor-specific DICOM attributes and variations in the encoding of radiotracer accumulation times.
Purpose: This technical note introduces a robust, vendor-specific SUV calculation strategy that addresses inconsistencies in current methodologies. We also integrate this strategy into an open-source software solution, Z-Rad, capable of converting raw PET DICOM data into body-weight normalized SUV NIfTI files.
Methods: Our SUV calculation strategy was developed by reviewing DICOM conformance statements from GE, Philips, and Siemens. Validation was conducted using real-world PET datasets, and the proposed strategy was compared to existing software solutions.
Results: Our SUV calculation approach demonstrated improved accuracy, particularly in resolving time-related discrepancies in the studied data. Our analysis also identified inconsistencies in the SUV calculation methods used by popular commercial and open-source software solutions, which do not fully account for vendor-specific DICOM attributes and PET image acquisition times. These limitations resulted in errors in SUV estimation reaching 33% when comparing our strategy to studied software.
Conclusions: The proposed vendor-specific SUV calculation strategy significantly enhances accuracy in PET imaging by addressing key inconsistencies caused by variations in DICOM attributes and image acquisition times across different vendors. This method effectively reduces SUV calculation errors and has been integrated into an open-source software, Z-Rad.
Keywords Standardized Uptake Value  $\cdot$ Positron Emission Tomography  $\cdot$ Z-Rad
1 Introduction
The standardized uptake value (SUV) is a fundamental metric in positron emission tomography (PET) 1, shown to be predictive for several tumor types 2, 3, essential for quantifying treatment response4, and is a precursor to more comprehensive metrics like total lesion glycolysis (TLG) or metabolic tumor volume (MTV) 5. However, the methods for calculating SUV from DICOM files vary across vendors 6, 7, 8, software solution9, and open-source datasets 10.
A significant contribution to the standardization of SUV calculations was made by Paul Kinahan et al. 11, who surveyed MIM, Philips, Siemens, and GE, and derived both vendor-specific 12 and vendor-neutral 13 SUV calculation strategies. Their vendor-specific approach 12 takes into account private DICOM attributes from different vendors, but does not completely account for the errors that occur when the Series Time (0008, 0031) is altered. In contrast, their vendor-neutral "happy path only" strategy 13 ignores all vendor-specific DICOM attributes and relies solely on Series Time (0008, 0031).
In this technical note, we present an adapted and comprehensive vendor-specific strategy for SUV calculation, addressing various pitfalls and challenges associated with real-world PET data. We provide a thorough analysis of existing SUV calculation methodologies, highlight potential issues, and propose a software solution to improve comparability of inter-vendor PET data.
2 Methods
2.1 SUV Definition
The general equation for the SUV normalized by the body weight ($SUV_{bw}$) is defined as follows:
$\text{SUV}_{bw}=\frac{A\cdot w}{D_{i}}\cdot 2^{\frac{t_{c}-t_{i}}{T_{1/2}}}\quad\left[\frac{g}{ml}\right],$ (1)
where $D_{i}$ is the injected dose of the radiotracer (in Bq), $w$ patient’s weight (in g), $A$ activity concentration (in Bq/mL), $t_{c}-t_{i}$ difference between measurement and radiotracer injection times respectively (in s), and $T_{1/2}$ radiotracer half-life time (in s).
2.2 Calculation of SUV from DICOM
In PET DICOM files, the Rescale Slope (0028,1053) and Rescale Intercept (0028,1052) are used to convert the pixel values stored in the image into physical units, as defined by Units (0054,1001). These units can represent measurements such as activity concentration in Becquerels per milliliter (BQML), grams per milliliter (GML), or counts (CNTS). This conversion is crucial because the raw pixel values directly from the scanner do not represent the radiotracer concentration. A useful parameter for further discussion, $A^{\prime}$, can be derived from the DICOM file using the following equation:
$A^{\prime}=P\cdot S+I,$ (2)
where $P$ represents the original stored pixel value, $S$ is the Rescale Slope (0028,1053), and $I$ is the Rescale Intercept (0028,1052). The parameter $A^{\prime}$ is measured in the units specified in Units attribute. If units are equal to BQML, then $A^{\prime}$ is equal to $A$ from Eq. 1.
Most discrepancies in SUV calculation from DICOM files originating from different vendors arise from different definitions of the measurement time $t_{c}$ (Eq. 1) and incorrect handling of different PET DICOM units. To address these issues, we derived a vendor-specific SUV calculation strategy based on the DICOM conformance statements from three major PET machine manufacturers: Philips 7, Siemens 8, and GE 6. Leveraging this information, we developed a unified framework to ensure consistency and accuracy in SUV definition across different systems (Fig. 1). A comprehensive list of all used DICOM attributes is provided in the Appendix.
2.2.1 Decay Correction Time
In metabolic PET imaging, the most common Units (0054, 1001) value is BQML. For this unit, Siemens, Philips, and GE agree that when the Decay Correction (0054, 1102) is set to ADMIN all images are decay-corrected to the tracer injection time $t_{i}$. This implies that $t_{c}$ is equal to $t_{i}$ in Eq. 1.
If the Decay Correction (0054, 1102) is START, vendors provide private DICOM attributes indicating the specific decay correction time $t_{c}$ to which the image is corrected. For instance, Siemens uses the private tag (0071, 1022), and GE uses (0009, 100d). If the corresponding tag is not present, or if the vendor is Philips - which does not have a dedicated private attribute - and the DICOM series includes multiple acquisition times, the decay correction time $t_{c}$ is defined as the earliest acquisition time in the series.
2.2.2 PET DICOM Units
Philips machines can also generate PET data with units of CNTS (counts). To calculate the SUV or the activity concentration (in Bq/mL) in this case, we refer to the equations provided in the Philips DICOM conformance statements 7:
$\text{SUV}=A^{\prime}\cdot f_{SUV},$ (3)
$A=A^{\prime}\cdot f_{BQML},$ (4)
where $A^{\prime}$ is defined by Equation 2, $f_{SUV}$ is the SUV Scale Factor (7053, 1000), and $f_{BQML}$ is the Activity Concentration Scale Factor (7053, 1009).
Another commonly used PET DICOM unit is GML (grams per milliliter). Although Siemens, GE, and Philips do not mention this unit in their DICOM conformance statements, the DICOM standard documentation 14 indicates that if the unit is GML and the SUV Type (0054, 1006) is set to BW (body-weight) or left unspecified, the SUV should be interpreted as a body-weight normalized and equal to $A^{\prime}$ (Eq. 2).
2.3 Validation datasets and Evaluated Software Solutions
To evaluate the proposed SUV calculation strategy (Fig. 1) and identify potential pitfalls, we tested it using PET datasets from our institution and from the autoPET challenge 10, which include metabolic PET data from Siemens, GE, and Philips (Table 1). Our strategy has been implemented in the open-source software Z-Rad 15, which can be used to convert raw PET DICOM files into body-weight normalized SUV NIfTI files. Additionally, we reviewed third-party software tools, such as the open-source software 3D Slicer and the commercial software MIM. Both solutions are widely used for reading PET DICOM files. This assessment helped us highlight any inconsistencies that may occur in their SUV calculations.
3 Results
A comparison of SUV values between Z-Rad, MIM, and 3D Slicer for GE data showed agreement among all three software for 59% of the images. For the remaining GE data, there was an average error of 31% for 3D Slicer and 32% for MIM resulting in underestimation of the SUV. This was caused by the image series time being shifted one hour earlier than the earliest acquisition time. In these cases, the GE private tag (0009, 100d) contained the correct decay correction time, which matched the earliest series acquisition time. MIM was able to detect when the series time is earlier than the injection time (16.5% of the data) and such images were not converted to SUV. However, MIM did not report errors when the series time was later than the injection time but still earlier than the earliest acquisition time (24.5% of the data). 3D Slicer filed to detect either of these two data related issues.
For Siemens data, the software solutions agreed for almost all images, except for three cases where a time difference of over an hour between the series and the earliest acquisition times resulted in an average SUV underestimation error of 33%. Another issue identified in the Siemens data was that the private Decay Correction Date Time attribute (0071, 1022) had the correct time but an incorrect date. Despite the series and acquisition dates being consistent, this could potentially lead to an erroneous addition of 86’400 seconds (24 hours) to $t_{c}$ (Eq. 1).
All Philips scans with BQML units showed a slight inconsistency between the series time and the earliest acquisition time. However, the maximum error in SUV estimation did not exceed 1.5%. For Philips images with CNTS, 3D Slicer’s PET module 9 did not apply the Philips-specific conversion factors (Eq. 3, Eq. 4), instead attempting to convert the data based on the Series Date and Series Time attributes. Similarly, MIM did not convert Philips PET DICOM files from CNTS to BQML, displaying the raw image in CNTS, which prevented any SUV-based clinical conclusions from being drawn. For the studied PET scans with CNTS, Philips provided private DICOM attributes that allow to convert image intensity to SUV (Eq. 3). However, since the SUV Type (0054,1006) was not specified in studied scans, we verified that Eq. 3 lead to SUV normalized to patient’s body weight in studied data.
4 Discussion
The vendor-neutral “happy path only” strategy proposed by Paul Kinahan et al. 13 assumes that $t_{c}$ is equal to the DICOM series time. However, this assumption is not always true. As shown in our results, this approach leads to errors when there is a mismatch between the earliest acquisition time and the series time, or when vendor-specific DICOM attributes are needed. The “happy path only” strategy is used in 3D Slicer, as stated in its documentation 9. While we do not know the specifics of SUV estimation methodology in MIM, our results suggest that it follows a similar “happy path only” vendor-neutral strategy, as it neither incorporated any vendor-specific DICOM attributes from the studied data nor correctly estimated $t_{c}$.
Another example of incorrect SUV estimation is the strategy used in the autoPET challenge 10. This approach relied on the acquisition time attribute of individual slices, resulting in an error that increased over the imaging duration. For DICOM data where the Series Time attribute remained unaltered, the error began at 0% when the acquisition time matched the series time but could grow to 21% for the last slices due to a 30-minute difference between the earliest and latest acquisition times. This led to images that failed to accurately depict SUV differences between body parts, as the error was not constant but gradually increased with each consecutive slice.
A closer examination of the vendor-specific strategy proposed by Paul Kinahan et al. 12 revealed that it addresses potential pitfalls in Philips CNTS data and accounts for GE and Siemens private tags, but only when the series time is later than the acquisition time. However, this approach does not apply to cases where the series time is earlier. As a result, this discrepancy leads to significant errors similar to those seen with the vendor-neutral "happy path only" strategy, resulting in average errors of up to 33% in the studied data.
5 Conclusions
The proposed vendor-specific SUV calculation strategy improves PET imaging accuracy by addressing key discrepancies in handling PET-related DICOM attributes across different vendors. Our method effectively mitigates errors caused by the misalignment of series time and acquisition time, issues commonly found in vendor-neutral approaches. Additionally, we have implemented this vendor-specific strategy in the open-source software Z-Rad, providing a ready-to-use tool for accurate SUV calculation from DICOM files of different PET scanner vendors.
6 Figures and Tables
Figure 1: Vendor-specific SUV calculation strategy.
Table 1: Vendor-specific SUV calculation strategy.
Data Source Vendor Manufacturer Model Name (0008,1090) Units (0054,1001) No. of PET Images
autoPET II 10 Siemens Biograph mCT, mCT Flow, Biograph 64 BQML (100%) 1014
USZ GE Discovery MI, Discovery 690, Discovery STE, Discovery 710, Discovery RX BQML (100%) 522
USZ Philips Allegro Body, GEMINI TF 64, GEMINI TF Big Bore, or not specified BQML(57%), CNTS (43%) 28
Appendix
Table 2: DICOM attributes
Attribute Name and Tag Vendor-specific Attribute Description
Rescale Intercept (0028, 1052) No The value b in relationship between stored values (P) and the output units. Units = m*P + b.
Rescale Slope (0028, 1053) No m in the equation specified by Rescale Intercept (0028, 1052).
Units (0054, 1001) No The units of the pixel values obtained after conversion from the stored pixel values (P) to pixel value units, as defined by Rescale Intercept (0028, 1052) and Rescale Slope (0028, 1053).
Corrected Image (0028, 0051) No One or more values that indicate which, if any, corrections have been applied to the images in this Series.
Manufacturer (0008, 0070) No Manufacturer of the equipment that produced the Composite Instances.
Decay Correction (0054, 1102) No The real-world event to which images in this Series were decay corrected.
Acquisition Time (0008, 0032) No The time the acquisition of data that resulted in this image started.
Series Time (0008, 0031) No Time the Series started.
SUV Type (0054, 1006) No Type of Standardized Uptake Value (SUV).
Patient’s Weight (0010, 1030) No Weight of the Patient, in kilograms.
Radionuclide Total Dose (0018, 1074) No The radiopharmaceutical dose administered to the patient measured in Becquerels (Bq) at the Radiopharmaceutical Start Time (0018,1072).
Radiotracer half-life time (0018, 1075) No The radionuclide half life, in seconds, that was used in the correction of this image.
Tracer injection time: (0018, 1072) No Time of start of administration. The actual time of radiopharmaceutical administration to the patient for imaging purposes, using the same time base as Series Time (0008, 0031).
SUV Scale Factor (7053, 1000) Philips This value only applies when Units (0054, 1001) is equal to CNTS. The SUV Scale Factor is used to convert the pixel data from counts to an SUV value. This is done by using the formula SUV Value = ((P * m) + b) * f, where: P - original stored pixel value, m - Rescale Slope (0028, 1053), b - Rescale Intercept (0028, 1052), f - SUV Scale Factor (7053, 1000). If the SUV Scale Factor is 0.0, then the pixel data cannot be converted from counts to an SUV value.
Activity Concentration Scale Factor (7053, 1009) Philips This value only applies when Units (0054, 1001) is equal to CNTS. The Activity Concentration Scale Factor is used to convert the pixel data from counts to Activity Concentration (in Bq/ml). This is done by using the formula Activity Concentration Value = ((P * m) + b) * f, where: P - original stored pixel value, m - Rescale Slope (0028, 1053), b - Rescale Intercept (0028, 1052), f - Activity Concentration Scale Factor (7053, 1009) If the Activity Concentration Scale Factor is 0.0, then the pixel data cannot be converted from counts to Activity Concentration.
Decay Correction DateTime (0071, 1022) Siemens -
Decay Correction DateTime (0009, 100d) GE -